# Phase-Space Topology in a Single-Atom Synthetic Dimension

### v2

@@ III Phase-space topology
which denotes the difference between phase-space Zak phase that illustrates the same quantum state geometry with the winding number in Eq. (6). The discrete change in $\Delta\gamma$ reflects the spin polarization of the SSH-like chain, illustrating the bulk–boundary correspondence in the Fock-state lattice. The winding number, defined along a single loop in $\phi$-space, captures the difference in the helical structure between $|n\rangle$ and $|n+1\rangle$, with the sign information encoded in the geometric parameters $R$ and $C$.
IV Quantum phase-space analysis
To interpret the observed topological properties in terms of quantum states, we introduce the quantum phase-space illustration of defect states. The $\phi$-dependent energy spectrum, as shown in Fig. 2 and Eq. (4), reflects the critical behavior of defect states emerging in phase space. Figure 5 illustrates the phase-space trajectories for $|v_{r}|-|v_{cr}|=\pm 0.75$, along with the Wigner functions of the zero-energy defect states at $\phi=\pi/4$. The squeezing axes are perpendicular to the black dotted line connecting the centers of the two defect states in different regions. As the difference $|v_{r}|-|v_{cr}|\to 0$, the defect states become infinitely squeezed along the squeezing axis, and both the center $C$ and the radius $R$ of the trajectory diverge. At $\phi=0$, the centers of the eigenstates in different phases coincide, as shown in Fig. 5, and the energy gap in Eq. (4) closes at the critical point $\Delta|v|=0$. In contrast, for $\phi\neq 0$, the squeezing axes remain parallel between the two regimes, but a spatial separation between the phase-space centers emerges and reaches a maximum at $\phi=\pi/2$. This separation diverges as $\Delta|v|\to 0$, resulting in a finite energy gap at the critical point. This critical behavior underlies the Dirac-cone structure shown in Eq. (4).
In contrast to classical systems, where topological phenomena emerge from the band structure, the topological features in our system are rooted in its entanglement structure and phase-space characteristics. The defect states possess topologically protected spin polarization, and the spin–boson entanglement in Eq. (2) induces a Zak phase in the single-atom quantum Rabi system. The phase-space geometry defines the corresponding topological invariant, the phase-space winding number, and the interplay between displacement and squeezing of the eigenstates leads to a conical energy spectrum. Our results highlight phase-space geometry as a powerful framework for realizing and probing topological phenomena in hybrid quantum spin–oscillator systems [56, 57], where topology arises from intrinsic quantum properties.
@@ IV Quantum phase-space analysis
Figure 5: Phase-space trajectories of $\alpha(\phi)$ for $w=2.5$. The red solid (blue dashed) trajectory corresponds to $|v_{r}|=1$, $|v_{cr}|=0.25$ ($|v_{r}|=0.25$, $|v_{cr}|=1$). The two trajectories intersect at $\phi=0$, and the Wigner functions of the zero-energy defect states at $\phi=\pi/4$ are shown for each regime. Gray dotted lines denote the squeezing axes, which are perpendicular to the black dot-dashed lines connecting the phase-space centers $\alpha(\phi)$ across the two regimes. While the squeezing axes remain parallel for all $\phi$, a spatial separation between the phase-space centers emerges and reaches its maximum at $\phi=\pi/2$. This separation diverges as $|v_{r}|-|v_{cr}|\to 0$ for $\phi\neq 0$, indicating the critical behavior that underlies the conical energy spectrum in Eq. (4).
V Discussion
@@
In this work, we have shown that a single-atom system described by the quantum Rabi model exhibits topological features characterized by its phase-space geometry. To overcome the absence of translational symmetry and the unconventional band structure of the Fock-state lattice, we introduced a phase-space winding number as a topological invariant, reflecting the quantized geometric phase. We showed that the proposed invariant describes the topologically protected fermionic spin state in the defect states of the Rabi SSH-like chain and links directly to the Zak phase of the Fock-state lattice. Furthermore, we analyzed the $\phi$-dependent dynamics in phase space, which leads to a conical energy spectrum at the critical point. Our results demonstrate topological features emerge from the intrinsic quantum nature of a single-atom synthetic lattice, establishing a compact and controllable framework for exploring topological physics. We demonstrate that topological phases, typically characterized by spatial symmetries in real-space lattices, can instead be identified through geometric features in the phase space of a synthetic Fock-state lattice.
The Hamiltonian in Eq. (1) can be implemented using trapped ions or superconducting circuits, where the coupling strengths $|v_{r}|$, $|v_{cr}|$, and $w$ correspond to sideband and carrier Rabi frequencies [19, 41, 42, 43, 44, 45]. In trapped-ion systems, typical red- and blue-sideband Rabi frequencies are on the order of tens of kHz, and $v_{r}$, $v_{cr}$, and $w$ can be independently tuned via radio-frequency (RF) amplitude modulation. We note, however, that the accessible parameter regime may be constrained by the Lamb–Dicke condition [41], which limits both the maximum achievable sideband coupling strength and the Fock-state population. In superconducting circuits, sideband Rabi frequencies can reach tens of MHz [45], allowing for faster dynamics. However, short coherence times [42] may restrict the evolution time, and the effective parameter regimes remain similar, as the topological features of the system are determined by relative coupling strengths rather than absolute values. In both platforms, the phase factors $\phi_{r}$ and $\phi_{cr}$ are tunable by controlling the relative phases of RF electronics used for sideband generation, enabling access to the desired phase-space geometry. These systems offer precise control over the drive parameters, making them suitable for the experimental demonstration of single-atom topological features via adiabatic preparation and manipulation of defect states.
An important implication of our results is the separation between the fermionic and bosonic nature of the defect state. While the spin polarization of the defect is topologically protected and robust against parameter noise, the bosonic properties, such as the occupation number $\langle n\rangle$ and the phase-space distribution, remain tunable via system parameters. This separation enables the realization of tunable, topologically protected quantum states, where the protected spin degree of freedom acts as a stable quantum memory and the bosonic mode can be exploited for control or interfacing. Such controllable yet protected quantum states may be useful in hybrid quantum computing architectures that require coherence in one degree of freedom and tunability in another [56, 57]. Furthermore, by utilizing multiple bosonic modes in hybrid quantum devices [58, 59, 60], our system can be extended to higher-dimensional synthetic lattices. These extended systems enable the exploration of novel topological phenomena arising from interactions between spin-protected defect modes. Such developments open up new possibilities for engineering topological matter, where topology originates from phase-space characteristics.
Acknowledgements.
VI Acknowledgments
@@
S. Yu was supported by Creative-Pioneering Researchers Program through Seoul National University. The rest of authors are supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant (No.RS-2022-II221040), and the National Research Foundation of Korea (NRF) grant (No. RS-2020-NR049232, No. RS-2024-00442855, No. RS-2024-00413957), all of which are funded by the Korean government (MSIT).

### v4

@@ Abstract
which defines the phase-space Zak phase that illustrates a quantum state geometry similar to the winding number in Eq. (6). The discrete change in $\Delta\gamma$ at $|v_{r}|=|v_{cr}|$ reflects the spin polarization of the SSH-like chain, illustrating the bulk–boundary correspondence in the FSL. The winding number, defined along a single loop in $\phi$-space, captures the difference in the helical structure between $|n\rangle$ and $|n+1\rangle$, with the sign information encoded in the geometric parameters $R$ and $C$.
@@
Figure 4: Phase-space trajectories of $\alpha(\phi)$ at $w=2.5$ in the frame rotated by $\phi_{cr}$. Red solid (blue dashed) trajectory corresponds to $|v_{r}|=1$, $|v_{cr}|=0.25$ ($|v_{r}|=0.25$, $|v_{cr}|=1$). The trajectories intersect at $\phi=0$, and the Wigner functions of the defect states at $\phi=\pi/4$ are shown for each regime. Gray dotted lines indicate the squeezing axes, perpendicular to the black dot–dashed lines connecting the phase-space centers $\alpha(\phi)$ between the two regimes.
Quantum phase-space analysis—To interpret the observed topological properties in terms of quantum states, we introduce the quantum phase-space illustration of defect states. The $\phi$-dependent energy spectrum, as shown in Eq. (4) and Appendix A, reflects the critical behavior of defect states emerging in phase space.
Figure 4 illustrates the phase-space trajectories of $\alpha(\phi)$ for $|v_{r}|-|v_{cr}|=\pm 0.75$, along with the Wigner functions $\mathcal{W}(\beta)\propto\langle\hat{D}(-\beta)\hat{\Pi}\hat{D}(\beta)\rangle$ of the zero-energy defect states at $\phi=\pi/4$, shown in the frame rotated by $\phi_{cr}$. Here, $\hat{\Pi}$ denotes the parity operator [66]. The squeezing axes are perpendicular to the black dot-dashed line connecting the centers of the two defect states in different regions. As the difference $|v_{r}|-|v_{cr}|\to 0$, the defect states become infinitely squeezed along the squeezing axis, and both the center $C$ and the radius $R$ of the trajectory diverge. At $\phi=0$, the centers of the eigenstates in different phases coincide, as shown in Fig. 4, and the energy gap in Eq. (4) closes at the critical point $\Delta|v|=0$. In contrast, for $\phi\neq 0$, the squeezing axes remain parallel between the two regimes, but a spatial separation between the phase-space centers emerges and reaches a maximum at $\phi=\pi/2$. This separation diverges as $\Delta|v|\to 0$, resulting in a finite energy gap at the critical point. This critical behavior underlies the Dirac-cone structure described in Eq. (4) and Appendix A.
In contrast to classical systems, where topological phenomena emerge from the band structure, the topological features in our system are rooted in its entanglement structure and phase-space characteristics. The defect state possesses topologically protected spin polarization, and the spin–boson entanglement in Eq. (2) induces a Zak phase in the single-atom quantum Rabi system. The phase-space geometry defines the corresponding topological invariant, the phase-space winding number, and the interplay between displacement and squeezing of the eigenstates leads to a conical energy spectrum. Our results highlight phase-space geometry as a powerful framework for realizing and probing topological phenomena in hybrid quantum spin–oscillator systems [67, 68], where topology arises from intrinsic quantum properties.
@@
Discussion—In this work, we have shown that a single-atom quantum Rabi system exhibits topological features characterized by its phase-space geometry. To address the absence of translational symmetry and the unconventional band structure in FSLs, we introduced a phase-space winding number as a topological invariant that reflects a quantized geometric phase. This invariant captures the topologically protected spin polarization of defect states in a semi-infinite QRM chain and connects directly to the Zak phase of the FSL. We further analyzed the $\phi$-dependent dynamics in phase space, which led to a conical energy spectrum emerging at the critical point. Our results show that topological features can arise from the intrinsic quantum nature of a single-atom synthetic lattice, providing a compact and controllable framework for exploring topological phenomena in quantum systems.
In particular, we showed that topological phases, typically characterized by spatial symmetries in real-space lattices, can instead be identified by geometric features in the phase space of a FSL. This provides a physically motivated, experimentally accessible topological quantity that bridges theoretical studies of semi-infinite FSLs with experimental realizations. For example, recent advances in bosonic state tomography [69, 70, 71] enable direct measurement of the phase-space winding number via tomographic reconstruction [51]. We briefly discuss experimental implementations on trapped-ion and superconducting-circuit platforms in Appendix D and in the Supplemental Material [51].
Another important implication of our results is the separation of spin and bosonic nature in the defect state. The spin polarization is topologically protected and robust to parameter noise, whereas bosonic properties, such as the occupation $\langle n\rangle$, designed selective flipping of spin polarization, and the phase-space distribution, remain tunable via system parameters. We demonstrated high-fidelity preparation of a nonclassical bosonic state under parameter noise and further enhanced its robustness via engineered dissipation to the $\left|g\right\rangle$ sublattice. These results open a pathway to the generation of precise nonclassical states even in the presence of noise, with potential applications in quantum metrology [72, 73, 74, 75, 76].
The Hamiltonian in Eq. (1) is widely used for gate operations in modern quantum processors [67], for example to generate spin–motion entanglement in trapped-ion systems [77]. The same interaction enables multiqubit entanglement via common motional modes [78, 79], and gates robust to parameter noise have been developed [80, 81, 82, 83, 84]. An alternative approach employs bosonic codes with spin-mediated interactions and readout [85, 71, 86, 87]. We investigated distinctive phase-space dynamics and the resulting conical spectrum induced by a coherent spin drive $w\hat{\sigma}_{x}$ at the critical point $|v_{r}|=|v_{cr}|$. These characteristics suggest potential routes to robust gate design and bosonic entangling operations. We expect our results to extend to topological phenomena arising from interactions between spin-protected defect modes, connecting quantum information science and topological physics through a phase-space perspective.
@@
Acknowledgments—Sunkyu Yu was supported by Creative-Pioneering Researchers Program through Seoul National University. The rest of authors are supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant (No.RS-2022-II221040), and the National Research Foundation of Korea (NRF) grant (No. RS-2020-NR049232, No. RS-2024-00442855, No. RS-2024-00413957), all of which are funded by the Korean government (MSIT).
Data availability—The simulation codes and data that support the findings of this article are openly available at [88]
Appendix
Figure 5: Energy spectrum $E^{\pm}_{n}$ of the Hamiltonian in Eq. (1) as a function of $|v_{r}|-|v_{cr}|$, with $w=5$. The spectrum is obtained via numerical diagonalization with a Fock-space cutoff $N_{\max}=3000$, for $\phi=0$ and $\phi=\pi/2$. The singular region $0.01\leq|v_{r}|-|v_{cr}|\leq 0.01$, which is excluded from the calculation, is masked out. Each line corresponds to an energy level $E_{n}$, and the color indicates its sign. Cross markers represent theoretical predictions from Eq. (2) for $1\leq n\leq 10$, showing good agreement with the numerical results, except near the singular region. When $\phi=\pi/2$, the spectrum displays a distinct feature near the point $|v_{r}|=|v_{cr}|$, as detailed in the main text.
Figure 6: Energy spectrum $E^{\pm}(x,\phi)$ given in Eq. (4) for $w=1$, $|v|=2$. A periodic array of conical Dirac points appears at $(q,\phi)=(0,m\pi)$ for $m\in\mathbb{Z}$.
Appendix A: Energy Spectrum—Figure 5 shows the energy spectra obtained via numerical diagonalization of the Hamiltonian in Eq. (1) for $\phi=0$ and $\phi=\pi/2$ under $|v_{r}|\neq|v_{cr}|$. The analytical expression in Eq. (2), indicated by cross markers, shows excellent agreement with the numerical results. Notably, the spectra exhibits qualitatively different behaviors near the critical point $|v_{r}|=|v_{cr}|$, depending on the values of $\phi$, as discussed in the main text. When $|v_{r}|=|v_{cr}|$, a conical energy spectrum emerges in a (1+1)(1\!+\!1)-dimensional space spanned by $x$ and $\phi$, as described by Eq. (4). Figure 6 illustrates the resulting periodic Dirac-cone structure in the $(x,\phi)$ space.
Figure 7: Phase-space trajectories with fixed $\phi_{r}$ and $\phi_{cr}$ swept over $0,2\pi]$. Other conditions are the same to Fig. 2. Red solid (blue dashed) curves correspond to the regime $|v_{r}|>|v_{cr}|$ ($|v_{r}|<|v_{cr}|$). The geometric properties remain the same, except the condition for whether the trajectory encircles the origin is reversed. Thus, the two topological phases can still be distinguished by the phase-space winding number.
Appendix B: Phase-Space Trajectory—All eigenstates in Eq. (2) are centered at
$\alpha_{r}(\phi_{r},\phi_{cr})=\frac{w(v_{cr}^{*}-v_{r})}{|v_{r}|-|v_{cr}|^{2}},$ (7)
where $\Re\{\alpha_{r}\}=\sqrt{2}\langle\hat{x}\rangle$ and $\Im\{\alpha_{r}\}=\sqrt{2}\langle\hat{p}\rangle$. Here, $\hat{x}$ and $\hat{p}$ are the position and momentum operators in the ladder basis of $\hat{a}$ and $\hat{a}^{\dagger}$, respectively. The additional degree of freedom $\phi_{cr}$ can be eliminated by moving to a rotated frame, $\alpha(\phi)=\alpha_{r}(\phi)e^{i\phi_{cr}}$. In this frame, the phase-space center becomes
$\alpha(\phi)=\frac{w(|v_{cr}|-|v_{r}|e^{i2\phi})}{|v_{r}|^{2}-|v_{cr}|^{2}}.$ (8)
Fixing $\phi_{cr}$ and varying $\phi_{r}$ from 0 to $2\pi$, the trajectories in Fig. 2 correspond to the positions of $\alpha(\phi)$ over $\phi\in[0,\pi]$ in the complex $\alpha$-plane. Since $\phi_{cr}$ is fixed, we have $\phi\propto-\theta$ for $2\theta=\phi_{cr}-\phi_{r}$, allowing the trajectories to be interpreted as geometric features of the eigenstates as the phase difference between the hopping terms $v_{r}$ and $v_{cr}$ varies from 0 to $2\pi$. A single loop of $\alpha(\phi)$ over $\phi\in[0,\pi]$ thus captures the helical structure of the eigenstates along the $\theta$-axis induced by $\hat{R}(\theta)$.
Alternatively, one may fix $\phi_{r}$ instead of $\phi_{cr}$. However, this choice does not affect the global property—whether the phase-space trajectories encircle the origin depends only on the relative values of $|v_{r}|$ and $|v_{cr}|$, as shown in Fig. 7. That is, the choice of fixed parameter determines only the sign of the relation between $\phi$ and $\theta$, while the geometric structure remains unchanged. Experimentally, the phase-space trajectories can be measured by preparing an eigenstate and tracking its phase-space distribution as a function of $\phi_{r}$ or $\phi_{cr}$, while keeping the other fixed. The choice of which phase parameter to vary depends solely on the experimentalist’s preference. The choice of phase parameter to vary is largely a matter of convenience and has no significant impact on the result.
The phase-space winding number can be calculated as
$W$ $=\frac{1}{2\pi i}\int_{0}^{\pi}d\phi\,\frac{d}{d\phi}\ln\alpha(\phi)=\frac{1}{2\pi i}\oint_{\mathcal{C}^{\prime}}\frac{1}{t}\,dt,$
where $t=|v_{cr}|-|v_{r}|e^{2i\phi}$. Here, $\mathcal{C}^{\prime}$ is the trajectory traced by $t$ as $\phi$ varies from 0 to $2\pi$. According to the residue theorem, the last integral evaluates to $2\pi i$ if the contour $\mathcal{C}^{\prime}$ encloses the origin, and to zero otherwise. Therefore, the winding number is given by
$2W=1-\mathrm{sgn}(|v_{cr}|-|v_{r}|).$
If instead we fix $\phi_{r}$ and define $\alpha(\phi)=\alpha_{r}(\phi)e^{-i\phi_{r}}$, the contour $\mathcal{C}^{\prime}$ becomes the trajectory traced by $t(\phi)=|v_{cr}|e^{-2i\phi}-|v_{r}|$, resulting in a winding number:
$2W=1+\mathrm{sgn}(|v_{cr}|-|v_{r}|).$
Figure 8: Fock-state populations of eigenstates at $w=v_{r}=1$ and $v_{cr}=0.25$. For clarity, the weaker hopping $v_{cr}$ is not shown. Solid and dashed lines indicate the stronger and weaker of the intra-cell $w$ and inter-cell $v_{r}$ couplings. Each population $|\langle n,A|\psi\rangle|^{2}$ for $A\in\{e,g\}$ is encoded by the opacity of the circles, with higher opacity indicating larger population. A domain wall appears at $n=1$, and the defect $\left|1,g\right\rangle$ is highlighted by a green box. The zero-energy defect state $\left|\psi_{0}\right\rangle$ is localized near the defect and resides on the $\left|g\right\rangle$ sublattice. In contrast, the populations of other eigenstates $\left|\psi_{k}^{+}\right\rangle$ are more spread over the FSL and occupy both sublattices.
Appendix C: Fock-state Populations— This section presents the Fock-state populations of the eigenstates $\left|\psi_{0}\right\rangle$ and $\left|\psi_{k}^{+}\right\rangle$ ($k\in\{1,2,3\}$) with parameters $v_{r}=w=1$ and $v_{cr}=0.25$ used in the numerical simulations of Fig. 3. In this case, a domain wall appears at site $n=1$, with the defect located at $\left|1,g\right\rangle$. Figure 8(a) shows the Fock-state population of the zero-energy defect state $\left|\psi_{0}\right\rangle$, which is localized near the defect, indicated by a green box. The defect state resides on the sublattice $\left|g\right\rangle$, and the topological protection of its spin polarization $\langle\psi_{0}|\hat{\sigma}_{z}|\psi_{0}\rangle$ prevents excitation into the other sublattice $\left|e\right\rangle$ under parameter noise. In contrast, the other eigenstates $\left|\psi_{k}^{+}\right\rangle$ exhibit Fock-state populations extended across the FSL, as shown in Fig. 8(b)–(d). These states occupy both sublattices and are therefore vulnerable to parameter noise, since their spin states are not protected.
Appendix D: Experimental Realization—The Hamiltonian in Eq. (1) can be implemented using trapped ions or superconducting circuits, where the coupling strengths $|v_{r}|$, $|v_{cr}|$, and $w$ correspond to sideband and carrier Rabi frequencies [19, 44, 45, 46, 47, 48]. Since the energy gap between $E_{0}$ and $E_{\pm 1}$ remains finite unless $|v_{r}|=|v_{cr}|$, the zero-energy defect state can be adiabatically prepared by first activating one of the sidebands ($v_{r}$ or $v_{cr}$) and then gradually increasing both $w$ and the remaining sideband strength. Geometric features, such as phase-space trajectories and distributions, can be estimated trough quantum state tomography [71, 69, 70]. The main limitation of this adiabatic protocol lies in the trade-off between the adiabatic timescale and the coherence time of the zero-energy state. The adiabatic timescale is determined by the energy gap $|E_{\pm 1}-E_{0}|=\sqrt{||v_{r}|-|v_{cr}||}$, and thus the coherence time must exceed this timescale to ensure high-fidelity state preparation and measurement.
In trapped-ion systems, typical red- and blue-sideband Rabi frequencies are on the order of tens of kHz, and $v_{r}$, $v_{cr}$, and $w$ can be independently tuned via radio-frequency (RF) amplitude modulation. The motional coherence time near the ground state can approach one second [95], making trapped-ion platforms promising candidates for probing phase-space characteristics through an adiabatic process. However, the coherence time is significantly shorter for highly excited Fock states, and the accessible parameter regime may be constrained by the Lamb–Dicke condition [44], which limits both the maximum achievable sideband coupling strength and the Fock-state population.
In superconducting circuits, sideband Rabi frequencies can reach tens of MHz [48], allowing for faster adiabatic simulations. However, the motional coherence time is relatively shorter than in trapped-ion systems, reaching up to tens of milliseconds [96, 97]. The effective parameter regimes remain similar to trapped-ion systems, as the system’s topological features are determined by relative coupling strengths rather than absolute values.
In both platforms, the phase factors $\phi_{r}$ and $\phi_{cr}$ are tunable by controlling the relative phases of RF electronics used for sideband generation, enabling precise control of the desired phase-space geometry. These systems offer precise control over the drive parameters, making them suitable for the experimental demonstration of single-atom topological features via adiabatic preparation and manipulation of defect states.